# Onsager reciprocal relations for charge and spin transport in periodically driven systems

### v1

@@ II.1 Charge off-diagonal dc conductivity
II.1.1 Nondriven systems
First, we consider the nondriven case with a magnetic field $\mathbf{H}$. In general, the Onsager reciprocal relation connects the transport coefficient for a certain transport phenomenon with another for the transport phenomenon obtained by applying the time-reversal operation to the original phenomenon Onsager1 ; Onsager2 ; Kubo . Therefore, the Onsager reciprocal relation for the charge off-diagonal dc conductivity with $\mathbf{H}$ is given by Onsager1 ; Onsager2 ; Kubo
@@ II.1.1 Nondriven systems
where $\sigma_{yx}^{\textrm{C}}(\mathbf{H})$ is the charge off-diagonal dc conductivity for the charge current $J_{\textrm{C}}^{y}$ generated perpendicular to the probe electric field $E_{x}$ in the presence of $\mathbf{H}$ [i.e., $J_{\textrm{C}}^{y}=\sigma_{yx}^{\textrm{C}}(\mathbf{H})E_{x}$], and $\sigma_{xy}^{\textrm{C}}(-\mathbf{H})$ is that for the charge current $J_{\textrm{C}}^{x}$ generated perpendicular to the probe electric field $E_{y}$ in the presence of $-\mathbf{H}$ [i.e., $J_{\textrm{C}}^{x}=\sigma_{xy}^{\textrm{C}}(-\mathbf{H})E_{y}$]. The factor $(-1)^{2}$ in Eq. (1) arises from the sign changes in the charge current operators appearing in the charge off-diagonal dc conductivity under the time-reversal operation [Fig. 3(a)]. Since $\sigma_{xy}^{\textrm{C}}(-\mathbf{H})=-\sigma_{xy}^{\textrm{C}}(\mathbf{H})$, the Onsager reciprocal relation can be reduced to
@@
This is often called the Onsager reciprocal relation, but it has been derived from a combination of the Onsager reciprocal relation and the additional symmetric property. This point is important to discuss whether a certain transport coefficient satisfies the Onsager reciprocal relation.
@@
In both cases, the charge off-diagonal dc conductivity has only the antisymmetric part $(\sigma_{yx}^{\textrm{C}}-\sigma_{xy}^{\textrm{C}})/2$ and thus can be regarded as the Hall conductivity, which is described by the antisymmetric part. Note that if there is no field breaking time-reversal symmetry, the Onsager reciprocal relation becomes $\sigma_{yx}^{\textrm{C}}=(-1)^{2}\sigma_{xy}^{\textrm{C}}=\sigma_{xy}^{\textrm{C}}$ Onsager1 , i.e., the charge off-diagonal dc conductivity has only the symmetric part $(\sigma_{yx}^{\textrm{C}}+\sigma_{xy}^{\textrm{C}})/2$.
II.1.2 Periodically driven systems
We now argue the Onsager reciprocal relations in periodically driven systems in two dimensions. To do this, we consider the nonequilibrium steady state under the pump field $\mathbf{A}_{\textrm{pump}}(t)$ with a time period $T_{\textrm{p}}$ [i.e., $\mathbf{A}_{\textrm{pump}}(t+T_{\textrm{p}})=\mathbf{A}_{\textrm{pump}}(t)$] and discuss the symmetric properties of the time-averaged charge off-diagonal dc conductivity [for its definition, see Eq. (51) for Q$=$C with Eq. (52)]. The following arguments can be extended to any other periodically driven systems.
@@ II.1.2 Periodically driven systems
Furthermore, the time-averaged charge off-diagonal dc conductivity changes its sign by switching the helicity of light NA-FloquetSHE ; NA-Onsager ,
@@
This property can be understood in terms of the symmetry of the charge current under the time-reversal operation, which switches the helicity of light NA-FloquetSHE . Combining Eqs. (7) and (8), we can reduce the Onsager reciprocal relation in this case to
@@
Therefore, the time-averaged charge off-diagonal dc conductivity with the pump field of CPL is restricted to be antisymmetric. This has been numerically confirmed in graphene driven by CPL NA-Onsager .
@@
Equation (10) has been also numerically confirmed in graphene driven by LPL NA-Onsager .
The difference between Eqs. (9) and (10) is due to the difference in time-reversal symmetry NA-Onsager . Note that CPL can break time-reversal symmetry Light-TrevBreaking , whereas LPL does not. We will check Eqs. (9) and (10) for periodically driven Sr2RuO4 in Sec. V B 1.
The situation becomes different in the presence of the pump field of BCPL BCPL-NatPhoto ; BCPL-Oka ; NA-BCPL , $\mathbf{A}_{\textrm{BCPL}}(t)=(A_{x}(t)\ A_{y}(t))^{T}$, where
@@
Since $\mathbf{A}_{\textrm{BCPL}}(-t)$ and $\mathbf{A}_{\textrm{BCPL}}(t)$ are not connected by a simple relation such as $\mathbf{A}_{\textrm{LCPL}}(-t)=\mathbf{A}_{\textrm{RCPL}}(t)$, we cannot rewrite Eq. (14) anymore in general. Therefore, in periodically driven systems with BCPL, the time-averaged charge off-diagonal dc conductivity is restricted to neither the antisymmetric nor symmetric part generally. We will validate this property numerically in Sec. V B 2.
II.2 Spin off-diagonal dc conductivity
We argue the Onsager reciprocal relations for the spin off-diagonal dc conductivity. We suppose that this conductivity is given by the correlation function between the charge current and spin current operators. This is valid in the linear-response regime of the probe field. Note that this conductivity is different from that for describing the spin current generated by the probe spin field, which is given by the correlation function between the spin current operators NA-MagnonTrans1 ; NA-MagnonTrans2 . We also suppose that the spin current remains unchanged under the time-reversal operation. Therefore, the following arguments are valid for any definition of the spin current as long as it is symmetric with respect to the time-reversal operation. This property against the time-reversal operation may be reasonable because the spin current is the flow of the spin angular momentum, such as $\mathbf{J}_{\textrm{S}}=\frac{\hbar}{2}(\mathbf{J}_{\uparrow}-\mathbf{J}_{\downarrow})$ NA-FloquetSHE , where $\mathbf{J}_{\sigma}$ is the current of spin-$\sigma$ electrons [Fig. 3(b)].
@@ II.2.1 Nondriven systems
where $\sigma_{yx}^{\textrm{S}}(\mathbf{H})$ is the spin off-diagonal dc conductivity for the spin current $J_{\textrm{S}}^{y}$ generated perpendicular to the probe electric field $E_{x}$ in the presence of $\mathbf{H}$ [i.e., $J_{\textrm{S}}^{y}=\sigma_{yx}^{\textrm{S}}(\mathbf{H})E_{x}$], and $\tilde{\sigma}_{xy}^{\textrm{S}}(-\mathbf{H})$ is another for the charge current $J_{\textrm{C}}^{x}$ generated perpendicular to the probe spin field $E_{y}^{\textrm{S}}$ (e.g., the gradient of the Zeeman field or of a spin-dependent chemical potential CorrectOnsager-JS1 ) in the presence of $-\mathbf{H}$ [i.e., $J_{\textrm{C}}^{x}=\tilde{\sigma}_{xy}^{\textrm{S}}(-\mathbf{H})E^{\textrm{S}}_{y}$]. The factor $(-1)$ in Eq. (17) arises from the properties that the spin off-diagonal dc conductivity is given by the correlation function between the charge current and spin current operators and that only the charge current operator changes its sign under the time-reversal operation (Fig. 3). This minus sign has been missing in some studies IncorrectOnsager-JS1 ; IncorrectOnsager-JS2 . Therefore, the violation of the Onsager reciprocal relation claimed in Ref. IncorrectOnsager-JS2, may be due to the incorrect treatment about the difference between time-reversal symmetries of the charge and spin currents. Then, Eq. (17) shows that the counterpart connected by the Onsager reciprocal relation for $\sigma_{yx}^{\textrm{S}}(\mathbf{H})$ is not $\sigma_{xy}^{\textrm{S}}(-\mathbf{H})$, but $\tilde{\sigma}_{xy}^{\textrm{S}}(-\mathbf{H})$, where $\sigma_{xy}^{\textrm{S}}(-\mathbf{H})$ is the spin off-diagonal dc conductivity for the spin current $J_{\textrm{S}}^{x}$ generated perpendicular to the probe electric field $E_{y}$ in the presence of $-\mathbf{H}$ [i.e., $J_{\textrm{S}}^{x}=\sigma_{xy}^{\textrm{S}}(-\mathbf{H})E_{y}$]. This is because the transport coefficient for a certain transport phenomenon is connected by the Onsager reciprocal relation with that for the transport phenomenon obtained by applying the time-reversal operation to the original phenomenon Onsager1 ; Onsager2 ; Kubo , as explained above.
@@
These results are consistent with some previous studies CorrectOnsager-JS1 ; CorrectOnsager-JS2 , although the origin of the minus sign appearing in the Onsager reciprocal relation is different.
@@
II.2.2 Periodically driven systems
We turn to the cases of the periodically driven systems in two dimensions. In the following, we discuss the symmetric properties of the time-averaged spin off-diagonal dc conductivity in the nonequilibrium steady state under $\mathbf{A}_{\textrm{pump}}(t)$. [For its definition, see Eq. (51) for Q$=$S with Eq. (52).] The arguments for any other periodically driven systems can be made in a similar way.
@@ II.2.2 Periodically driven systems
In addition, the time-averaged spin off-diagonal dc conductivity is independent of the helicity of light NA-FloquetSHE ,
@@
This equation can be understood in terms of the symmetry of the spin current under the time-reversal operation NA-FloquetSHE . Combining Eq. (22) with Eqs. (20) and (21), we obtain the Onsager reciprocal relation in this case,
@@
Therefore, the time-averaged spin off-diagonal dc conductivity is restricted to be antisymmetric in the periodically driven system with the pump field of CPL or LPL. We will check this property numerically in Sec. V A 1.
@@
As well as Eq. (14), this equation cannot be rewritten anymore in general. However, as we will show in Sec. V A 2, the antisymmetric part dominates the time-averaged spin off-diagonal dc conductivity in the nonequilibrium steady states with BCPL. This is different from the result of the time-averaged charge off-diagonal dc conductivity with BCPL (see Sec. V B 2).
III Model
Figure 5: (a) The nearest neighbor and next nearest neighbor hopping integrals for the $t_{2g}$-orbital electrons of a Ru ion on the square lattice. The $d_{yz}$, $d_{zx}$, and $d_{xy}$ correspond to the $d_{yz}$, $d_{zx}$, and $d_{xy}$ orbitals, respectively. (b) The band structure and (c) Fermi surface of our model for nondriven Sr2RuO4 at $T_{\textrm{b}}=0.02$ eV for a $N_{x}\times N_{y}$ mesh with $N_{x}=N_{y}=100$. In panel (b), the energies are measured from the chemical potential $\mu$.
To show the validity of our general arguments made in Sec. II, we consider a concrete model and analyze its time-averaged spin and charge off-diagonal dc conductivities. Sections III, IV, and V are devoted to this analysis.
In this section, we introduce the concrete model for periodically driven electron systems. As the concrete model, we consider a periodically driven multiorbital metal with weak coupling to a heat bath. Our model Hamiltonian consists of three parts NA-FloquetSHE :
@@ III Model
where $H_{\textrm{s}}(t)$ is the system Hamiltonian, $H_{\textrm{b}}$ is the bath Hamiltonian, and $H_{\textrm{sb}}$ is the system-bath coupling Hamiltonian. As $H_{\textrm{s}}(t)$, we have considered the periodically driven Sr2RuO4 because it is suitable for achieving the finite time-averaged spin and charge off-diagonal dc conductivities NA-FloquetSHE . In addition to $H_{\textrm{s}}(t)$, we have considered $H_{\textrm{b}}$ and $H_{\textrm{sb}}$ because we suppose that our periodically driven system can reach a nonequilibrium steady state due to the coupling with the heat bath Tsuji ; Mikami ; NA-FloquetSHE . If such a relaxation mechanism is absent, a periodically driven system reaches an infinite-temperature state due to the heating of the pump field Heating1 ; Heating2 . Hereafter, we use the unit $\hbar=k_{\textrm{B}}=c=a_{\textrm{NN}}=1$, where $a_{\textrm{NN}}$ is the distance between nearest-neighbor sites on a square lattice.
III.1 System Hamiltonian $H_{\textrm{s}}(t)$
@@
We choose the parameters of $H_{\textrm{s}}(t)$ to reproduce the electronic structure of Sr2RuO4 near the Fermi level. The hopping integrals $t_{ij}^{ab}$’s in Eq. (28) are parametrized by $t_{1}$, $t_{2}$, $t_{3}$, $t_{4}$, and $t_{5}$ NA-FloquetSHE ; NA-FLEX ; NA-Ru-model ; the first three ones are nearest-neighbor terms, and the others are next-nearest-neighbor ones [Fig. 5(a)]. We choose their values as follows: $(t_{1},t_{2},t_{3},t_{4},t_{5})=(0.675,0.09,0.45,0.18,0.03)$ (eV) NA-Ru-model . We also set $\xi=0.17$ eV SOC-Oguchi . Then, we determine $\mu$ from the condition that the electron number per site is four; the value of $\mu$ is fixed at that determined for $\mathbf{A}(t)=\mathbf{0}$ in the analyses shown in Sec. V. Figures 5(b) and 5(c) show the band structure and Fermi surface of our model for Sr2RuO4 with $\mathbf{A}(t)=\mathbf{0}$. The obtained Fermi surface is consistent with that observed experimentally in nondriven Sr2RuO4 Sr2RuO4-ARPES . Note that Sr2RuO4 is a material with the simple electronic structure in which the spin off-diagonal dc conductivity is finite Kon-SHE-Ru ; Mizo .
III.2 Bath Hamiltonian $H_{\textrm{b}}$
$H_{\textrm{b}}$ is the Hamiltonian for the Büttiker-type heat bath HeatBath1 ; HeatBath2 ; Tsuji ; Mikami ; NA-FloquetSHE :

### v2

@@ 2.1 Charge off-diagonal dc conductivity
2.1.1 Nondriven systems
First, we consider the nondriven case with a magnetic field $\mathbf{H}$. In general, the Onsager reciprocal relation connects the transport coefficient for a certain transport phenomenon with another for the transport phenomenon obtained by applying the time-reversal operation to the original phenomenon [19, 20, 21]. Therefore, the Onsager reciprocal relation for the charge off-diagonal dc conductivity with $\mathbf{H}$ is given by [19, 20, 21]
@@ 2.1.1 Nondriven systems
where $\sigma_{yx}^{\textrm{C}}(\mathbf{H})$ is the charge off-diagonal dc conductivity for the charge current $J_{\textrm{C}}^{y}$ generated perpendicular to the probe electric field $E_{x}$ in the presence of $\mathbf{H}$ [i.e., $J_{\textrm{C}}^{y}=\sigma_{yx}^{\textrm{C}}(\mathbf{H})E_{x}$], and $\sigma_{xy}^{\textrm{C}}(-\mathbf{H})$ is that for the charge current $J_{\textrm{C}}^{x}$ generated perpendicular to the probe electric field $E_{y}$ in the presence of $\mathbf{H}$ [i.e., $J_{\textrm{C}}^{x}=\sigma_{xy}^{\textrm{C}}(-\mathbf{H})E_{y}$]. The factor $(-1)^{2}$ in Eq. (1) arises from the sign changes in the charge current operators appearing in the charge off-diagonal dc conductivity under the time-reversal operation [Fig. 3(a)]. Then, we suppose that the charge off-diagonal dc conductivity satisfies $\sigma_{xy}^{\textrm{C}}(-\mathbf{H})=-\sigma_{xy}^{\textrm{C}}(\mathbf{H})$, which usually holds in the case with $\mathbf{H}=(0\ 0\ H)^{T}$. By combining this relation with Eq. (1), the Onsager reciprocal relation can be reduced to
@@
This is often called the Onsager reciprocal relation, but it has been derived from a combination of the Onsager reciprocal relation Eq. (1) and the additional symmetric property [i.e., $\sigma_{xy}^{\textrm{C}}(-\mathbf{H})=-\sigma_{xy}^{\textrm{C}}(\mathbf{H})$]. This point is important to discuss whether a certain transport coefficient satisfies the Onsager reciprocal relation. If there is no such additional symmetric property, whether the Onsager reciprocal relation holds should be discussed using the equation such as Eq. (1); this is true in the case with BCPL, as we will show in Sect. 2.1.2.
@@
In driving this equation, we have supposed that $\sigma_{xy}^{\textrm{C}}(-\mathbf{M})=-\sigma_{xy}^{\textrm{C}}(\mathbf{M})$ is satisfied.
In both cases, the charge off-diagonal dc conductivity has only the antisymmetric part $(\sigma_{yx}^{\textrm{C}}-\sigma_{xy}^{\textrm{C}})/2$ and thus can be regarded as the Hall conductivity, which is described by the antisymmetric part. This antisymmetric part can be finite only with broken time-reversal symmetry [19].
2.1.2 Periodically driven systems
We now argue the Onsager reciprocal relations in periodically driven systems in two dimensions. To do this, we consider the nonequilibrium steady state under the pump field $\mathbf{A}_{\textrm{pump}}(t)$ with a time period $T_{\textrm{p}}$ [i.e., $\mathbf{A}_{\textrm{pump}}(t+T_{\textrm{p}})=\mathbf{A}_{\textrm{pump}}(t)$] and discuss the symmetric properties of the time-averaged charge off-diagonal dc conductivity [for its definition, see Eq. (51) for Q==C with Eq. (52)]. The following arguments can be extended to any other periodically driven systems.
@@ 2.1.2 Periodically driven systems
Furthermore, we suppose that the time-averaged charge off-diagonal dc conductivity changes its sign by switching the helicity of light [11, 27],
@@
This property can be understood in terms of the symmetry of the charge current under the time-reversal operation, which switches the helicity of light [11]. Combining Eqs. (7) and (8), we can reduce the Onsager reciprocal relation in this case to
@@
Therefore, the time-averaged charge off-diagonal dc conductivity with the pump field of CPL is restricted to be antisymmetric. This has been numerically confirmed in graphene driven by CPL [27].
@@
Equation (10) has been also numerically confirmed in graphene driven by LPL [27].
The difference between Eqs. (9) and (10) is due to the difference in time-reversal symmetry [27]. Note that CPL can break time-reversal symmetry [31], whereas LPL does not. We will check Eqs. (9) and (10) for periodically driven Sr2RuO4 in Sect. 5.2.1.
The situation becomes different in the presence of the pump field of BCPL [28, 29, 30], $\mathbf{A}_{\textrm{BCPL}}(t)=(A_{x}(t)\ A_{y}(t))^{T}$, where
@@
Since $\mathbf{A}_{\textrm{BCPL}}(-t)$ and $\mathbf{A}_{\textrm{BCPL}}(t)$ are not connected by a simple relation such as $\mathbf{A}_{\textrm{LCPL}}(-t)=\mathbf{A}_{\textrm{RCPL}}(t)$, we cannot rewrite Eq. (14) anymore in general. Therefore, in periodically driven systems with BCPL, the time-averaged charge off-diagonal dc conductivity is restricted to neither the antisymmetric nor symmetric part generally. We will validate this property numerically in Sect. 5.2.2.
2.2 Spin off-diagonal dc conductivity
We argue the Onsager reciprocal relations for the spin off-diagonal dc conductivity. We suppose that this conductivity is given by the correlation function between the charge current and spin current operators. This is valid in the linear-response regime of the probe field. Note that this conductivity is different from that for describing the spin current generated by the probe spin field, which is given by the correlation function between the spin current operators [32, 33]. We also suppose that the spin current remains unchanged under the time-reversal operation. Therefore, the following arguments are valid for any definition of the spin current as long as it is symmetric with respect to the time-reversal operation. This property against the time-reversal operation may be reasonable because the spin current is the flow of the spin angular momentum, such as $\mathbf{J}_{\textrm{S}}=\frac{\hbar}{2}(\mathbf{J}_{\uparrow}-\mathbf{J}_{\downarrow})$ [11], where $\mathbf{J}_{\sigma}$ is the current of spin-$\sigma$ electrons [Fig. 3(b)].
@@ 2.2.1 Nondriven systems
where $\sigma_{yx}^{\textrm{S}}(\mathbf{H})$ is the spin off-diagonal dc conductivity for the spin current $J_{\textrm{S}}^{y}$ generated perpendicular to the probe electric field $E_{x}$ in the presence of $\mathbf{H}$ [i.e., $J_{\textrm{S}}^{y}=\sigma_{yx}^{\textrm{S}}(\mathbf{H})E_{x}$], and $\tilde{\sigma}_{xy}^{\textrm{S}}(-\mathbf{H})$ is another for the charge current $J_{\textrm{C}}^{x}$ generated perpendicular to the probe spin field $E_{y}^{\textrm{S}}$ (e.g., the gradient of the Zeeman field or of a spin-dependent chemical potential [34]) in the presence of $\mathbf{H}$ [i.e., $J_{\textrm{C}}^{x}=\tilde{\sigma}_{xy}^{\textrm{S}}(-\mathbf{H})E^{\textrm{S}}_{y}$]. The factor (1)(-1) in Eq. (17) arises from the properties that the spin off-diagonal dc conductivity is given by the correlation function between the charge current and spin current operators and that only the charge current operator changes its sign under the time-reversal operation (Fig. 3). This minus sign has been missing in some studies [35, 36]. Therefore, the violation of the Onsager reciprocal relation claimed in Ref. [36] may be due to the incorrect treatment about the difference between time-reversal symmetries of the charge and spin currents. Then, Eq. (17) shows that the counterpart connected by the Onsager reciprocal relation for $\sigma_{yx}^{\textrm{S}}(\mathbf{H})$ is not $\sigma_{xy}^{\textrm{S}}(-\mathbf{H})$, but $\tilde{\sigma}_{xy}^{\textrm{S}}(-\mathbf{H})$, where $\sigma_{xy}^{\textrm{S}}(-\mathbf{H})$ is the spin off-diagonal dc conductivity for the spin current $J_{\textrm{S}}^{x}$ generated perpendicular to the probe electric field $E_{y}$ in the presence of $\mathbf{H}$ [i.e., $J_{\textrm{S}}^{x}=\sigma_{xy}^{\textrm{S}}(-\mathbf{H})E_{y}$]. This is because the transport coefficient for a certain transport phenomenon is connected by the Onsager reciprocal relation with that for the transport phenomenon obtained by applying the time-reversal operation to the original phenomenon [19, 20, 21], as explained above.
@@
These results are consistent with some previous studies [34, 37], although the origin of the minus sign appearing in the Onsager reciprocal relation is different.
@@
2.2.2 Periodically driven systems
We turn to the cases of the periodically driven systems in two dimensions. In the following, we discuss the symmetric properties of the time-averaged spin off-diagonal dc conductivity in the nonequilibrium steady state under $\mathbf{A}_{\textrm{pump}}(t)$. [For its definition, see Eq. (51) for Q==S with Eq. (52).] The arguments for any other periodically driven systems can be made in a similar way.
@@ 2.2.2 Periodically driven systems
In addition, the time-averaged spin off-diagonal dc conductivity is independent of the helicity of light [11],
@@
This equation can be understood in terms of the symmetry of the spin current under the time-reversal operation [11]. Combining Eq. (22) with Eqs. (20) and (21), we obtain the Onsager reciprocal relation in this case,
@@
Therefore, the time-averaged spin off-diagonal dc conductivity is restricted to be antisymmetric in the periodically driven system with the pump field of CPL or LPL. We will check this property numerically in Sect. 5.1.1.
@@
As well as Eq. (14), this equation cannot be rewritten anymore in general. However, as we will show in Sect. 5.1.2, the antisymmetric part dominates the time-averaged spin off-diagonal dc conductivity in the nonequilibrium steady states with BCPL. This is different from the result of the time-averaged charge off-diagonal dc conductivity with BCPL (see Sect. 5.2.2).
3 Model
Figure 5: (Color online) (a) The nearest neighbor and next nearest neighbor hopping integrals for the $t_{2g}$-orbital electrons of Ru ions on the square lattice. The $d_{yz}$, $d_{zx}$, and $d_{xy}$ correspond to the $d_{yz}$, $d_{zx}$, and $d_{xy}$ orbitals, respectively. (b) The band structure and (c) Fermi surface of our model for nondriven Sr2RuO4 at $T_{\textrm{b}}=0.02$ eV for a $N_{x}\times N_{y}$ mesh with $N_{x}=N_{y}=100$. In panel (b), the energies are measured from the chemical potential $\mu$.
To show the validity of our general arguments made in Sect. 2, we consider a concrete model and analyze its time-averaged spin and charge off-diagonal dc conductivities. Sections 3, 4, and 5 are devoted to this analysis.
In this section, we introduce the concrete model for periodically driven electron systems. As the concrete model, we consider a periodically driven multiorbital metal with weak coupling to a heat bath. Our model Hamiltonian consists of three parts [11]:
@@ 3 Model
where $H_{\textrm{s}}(t)$ is the system Hamiltonian, $H_{\textrm{b}}$ is the bath Hamiltonian, and $H_{\textrm{sb}}$ is the system-bath coupling Hamiltonian. As $H_{\textrm{s}}(t)$, we have considered the periodically driven Sr2RuO4 because it is suitable for achieving the finite time-averaged spin and charge off-diagonal dc conductivities [11]. In addition to $H_{\textrm{s}}(t)$, we have considered $H_{\textrm{b}}$ and $H_{\textrm{sb}}$ because we suppose that our periodically driven system can reach a nonequilibrium steady state due to the coupling with the heat bath [38, 8, 11]. If such a relaxation mechanism is absent, a periodically driven system reaches an infinite-temperature state due to the heating of the pump field [39, 40]. Hereafter, we use the unit $\hbar=k_{\textrm{B}}=c=a_{\textrm{NN}}=1$, where $a_{\textrm{NN}}$ is the distance between nearest-neighbor sites on a square lattice.
3.1 System Hamiltonian $H_{\textrm{s}}(t)$
@@
We choose the parameters of $H_{\textrm{s}}(t)$ to reproduce the electronic structure of Sr2RuO4 near the Fermi level. The hopping integrals $t_{ij}^{ab}$’s in Eq. (28) are parametrized by $t_{1}$, $t_{2}$, $t_{3}$, $t_{4}$, and $t_{5}$ [11, 42, 41]; the first three ones are nearest-neighbor terms, and the others are next-nearest-neighbor ones [Fig. 5(a)]. We choose their values as follows: $(t_{1},t_{2},t_{3},t_{4},t_{5})=(0.675,0.09,0.45,0.18,0.03)$ (eV) [41]. We also set $\xi=0.17$ eV [43]. Then, we determine $\mu$ from the condition that the electron number per site is four; the value of $\mu$ is fixed at that determined for $\mathbf{A}(t)=\mathbf{0}$ in the analyses shown in Sect. 5. Figures 5(b) and 5(c) show the band structure and Fermi surface of our model for Sr2RuO4 with $\mathbf{A}(t)=\mathbf{0}$. The obtained Fermi surface is consistent with that observed experimentally in nondriven Sr2RuO4 [44]. Note that Sr2RuO4 is a material with the simple electronic structure in which the spin off-diagonal dc conductivity is finite [45, 46].
3.2 Bath Hamiltonian $H_{\textrm{b}}$
$H_{\textrm{b}}$ is the Hamiltonian for the Büttiker-type heat bath [47, 48, 38, 8, 11]: